OpenStudy (anonymous):

Solve 10 ^(2x+9)=290 Express in terms of natura logarithms

OpenStudy (anonymous):

x=[log290/2log10-9/2]

OpenStudy (wolf1728):

10^(2x+9) = 290 (2x+9) * ln(10) = ln(290) (2x+9) * 2.302585093 = 5.669880923 4.605170186x + 20.723265837 = 5.669880923 4.605170186x = -15.053384914 I think you can solve for x. By the way, even though you wanted this solved by natural logs, base 10 logs or any other base would yield the same answer.